OpenStudy (anonymous):

Find tan(π/6) Find cot(π/3). Find tan(π/4). Find sec(π/4). All without using a calculator.

OpenStudy (perl):

hi sfb

OpenStudy (perl):

sorry i lost my temper earlier

OpenStudy (anonymous):

It's fine, @perl. Sorry for being a fluttering idiot.

OpenStudy (perl):

i was wrong to say that

OpenStudy (perl):

thanks for your understanding :D

OpenStudy (anonymous):

\[\tan \theta =\frac{ opposite }{ adjacent } \]\[\cot \theta = \frac{ 1 }{ \tan \theta }\]\[\sec \theta = \frac{ 1 }{ \cos \theta }\]

OpenStudy (campbell_st):

this question is asking about exact value trig ratios you need the triangles shown below  so find the ratios by looking at the diagrams for cot(pi/3) find tan(pi/3) then take the reciprocal of the fraction (flip it) and for sec (pi/4) find cos(pi/4) and flip the fraction for sec